# Cross-View Geo-Localization with Street-View and VHR Satellite Imagery in Decentrality Settings

### v1

@@ Abstract
Cross-View Geo-Localization tackles the problem of image geo-localization in GNSS-denied environments by matching street-view query images with geo-tagged aerial-view reference images. However, existing datasets and methods often assume center-aligned settings or only consider limited decentrality (i.e., the offset of the query image from the reference image center). This assumption overlooks the challenges present in real-world applications, where large decentrality can significantly enhance localization efficiency but simultaneously lead to a substantial degradation in localization accuracy. To address this limitation, we introduce CVSat, a novel dataset designed to evaluate cross-view geo-localization with a large geographic scope and diverse landscapes, emphasizing the decentrality issue. Meanwhile, we propose AuxGeo (Auxiliary Enhanced Geo-Localization), which leverages a multi-metric optimization strategy with two novel modules: the Bird’s-eye view Intermediary Module (BIM) and the Position Constraint Module (PCM). BIM uses bird’s-eye view images derived from street-view panoramas as an intermediary, simplifying the “cross-view challenge with decentrality” to a cross-view problem and a decentrality problem. PCM leverages position priors between cross-view images to establish multi-grained alignment constraints. These modules improve the performance of cross-view geo-localization with the decentrality problem. Extensive experiments demonstrate that AuxGeo outperforms previous methods on our proposed CVSat dataset, mitigating the issue of large decentrality, and also achieves state-of-the-art performance on existing public datasets such as CVUSA, CVACT, and VIGOR. The code and dataset will be made available at https://github.com/SummerpanKing/CVSat.
@@
Geo-localization , Image retrieval , Cross-view data fusion , Urban perception , Representation learning
@@
[label1]organization=School of Remote Sensing and Information Engineering, Wuhan University,city=Wuhan, postcode=430079, state=Hubei, country=China
[…]
[label2]organization=Ant Group,city=Hangzhou, postcode=310023, state=Zhejiang, country=China
[…]
[label3]organization=Technology Innovation Center for Collaborative Applications of Natural Resources Data in GBA, Ministry of Natural Resources,city=Guangzhou, postcode=510075, state=Guangdong, country=China
@@ 1 Introduction
Cross-View Geo-Localization (CVGL) tackles the problem of image geo-localization in GNSS-denied environments, by matching street-view query images with aerial-view geo-tagged very high resolution (VHR) satellite reference images from the database. It is particularly beneficial in complex environments where GNSS signals are unavailable, jammed, or compromised, such as urban canyons with dense building clusters or heavily forested areas [1]. This technology offers us an alternative way to localize ourselves in real scenarios, enabling emerging applications such as autonomous driving [2, 3, 4], robotic navigation [5], augmented reality [6], and disaster response [7].
@@
CVGL is challenging primarily due to the significant appearance changes between street-view and aerial-view images caused by viewpoint variations [8, 9]. Moreover, in real-world applications, it is not likely to have perfectly matched (center-aligned) reference images available for each query image, since the reference database is established with features generated by aerial images in advance. CVGL has to be robust and able to tolerate offsets between a street-view query and its corresponding aerial-view reference. Thus, we introduce the concept of decentrality, measuring the offset level between a query image position w.r.t. the center of its reference image. For the same search region, large decentrality requires less overlap between reference images, significantly reducing the number of reference images in the database and thereby improving the retrieval efficiency, which is essential for localization tasks. However, large decentrality presents substantial challenges to CVGL by diminishing content similarity between cross-view images and introducing disruptions caused by non-co-visible elements. Addressing these challenges requires robust partial matching mechanisms to effectively align cross-view images, as demonstrated in Figure 1. A detailed statistical analysis included in section 3.2 also highlights the importance of the decentrality issue in real-world applications.
@@
In datasets like CVUSA [10] and CVACT [11], the correspondences between cross-view images are assumed to be well center-aligned, which is over-simplified for real-world applications. The VIGOR dataset [12] relaxes this restriction of center-aligned correspondence to some extent, while, a query image is still located within a quite small area around the center of its reference image, see Figure 2(a). While the VIGOR dataset defines some ”semi-positive” street-view images with larger decentrality, there is always an alternative reference image available to form a ”positive” pair for every street-view image. As a result, the VIGOR dataset is not well-suited for effectively investigating the decentrality issue, even if its original structure is reorganized. Therefore, we propose a novel dataset, CVSat, to evaluate CVGL in more realistic settings. CVSat distinguishes itself from previous datasets in two key ways. First, it allows for the evaluation of CVGL with larger decentrality (see Figure 2(a)) by sampling reference images seamlessly with a low overlap (12.5%). Second, CVSat comprises cross-view images collected from 8 cities worldwide, covering both urban and suburban areas with diverse landscapes, where the area of each reference image exceeds 400 $km^{2}$. This new large dataset enables a step forward in studying CVGL in a practical setting. A detailed comparison of decentrality among the above datasets is included in section 3.3.
@@
Large decentrality presents significant challenges for existing methods. Methods using polar transform, which convert aerial-view images into street-view-like images [13], are unsuitable without center-aligned settings. Other state-of-the-art methods [14, 15] do not take into consideration the impact of offsets between street-view and aerial-view images across a broader portion of decentrality, resulting in a significant decrease in retrieval precision. Therefore, we propose a novel method, AuxGeo (auxiliary enhanced geolocalization), to address the aforementioned problems. The AuxGeo employs multi-metric optimization with two innovative modules incorporated as auxiliary tasks to enhance the representation ability of the backbone. To mitigate viewpoint variation involving decentrality, we introduce the Bird’s-eye view Intermediary Module (BIM). The BIM uses BEV images derived from street-view panoramas as intermediaries, establishing additional connections between street-BEV and aerial-BEV images, simplifying the “cross-view challenge with decentrality” to a cross-view problem and a decentrality problem. To mitigate the problem of decentrality, we introduce the Position Constraint Module (PCM). The PCM leverages prior position correspondences between street-view and aerial-view images as supervision, establishing alignments in a multi-grained manner and enhancing the backbone’s ability to learn cross-view correspondences for feature representation even with significant decentrality.
@@
Extensive experiments demonstrate that AuxGeo improves retrieval precision, surpassing previous methods on our proposed CVSat dataset and existing public datasets. Our main contributions can be summarized as follows:
  • 1.
@@
  • 2.
    We propose CVSat, the first public dataset designed to comprehensively evaluate CVGL methods in decentrality settings. It includes cross-view images from diverse geographic locations with extensive land coverage.
  • 3.
@@
  • 4.
    Our proposed method AuxGeo outperforms previous approaches on the CVSat dataset, showing progressively greater improvements over the current methods as decentrality increases. It also achieves state-of-the-art performance on existing public datasets. Importantly, AuxGeo incurs no additional pre-processes or computational costs during inference.
The remainder of this paper is structured as follows: Section 2 provides a comprehensive review of related work, including existing datasets and geo-localization methods in the cross-view topic. Section 3 introduces the proposed CVSat dataset in detail and offers an in-depth analysis of the importance the decentrality issue. Section 4 outlines the proposed method, AuxGeo. Experimental results and discussions are presented in Sections 5 and 6, respectively. Finally, Section 7 concludes the paper with an analysis of the method’s advantages and limitations.
@@ 2.1 Cross-View Datasets
There are many datasets that have been introduced for cross-view geo-localization in the past decade [16, 17, 18]. Among them, the most widely used datasets are CVUSA, CVACT, VIGOR and University-1652.
The original CVUSA is a huge dataset containing over 1 million ground and aerial images from multiple cities in the United States, first proposed by Workman et al. [19]. The current used version of CVUSA is a subset made by Zhai et al. [10], including 35,532 street-aerial image pairs for training and 8,884 pairs for testing. Similar to CVUSA, Liu et al. [11] proposed CVACT which has 35,532 image pairs for training, 8,884 pairs for validation and 92,802 pairs for testing. Extending beyond conventional one-to-one retrieval, Zhu et al. [12] introduced VIGOR. This dataset comprises 105,214 street-view images and 90,618 aerial-view images, uniquely characterized by its assumption of random placements within the target area without center-aligned settings. VIGOR is regarded as the most realistic dataset for cross-view geo-localization, as it extends the relationships between cross-view images beyond merely center-aligned correspondence. However, even with this advancement, the decentrality of positive samples in VIGOR is still confined to a relatively centered area, which restricts its potential for a more thorough exploration of the issue. University-1652 [20] is the first dataset to simultaneously include images from satellite, synthetic drone, and ground levels, providing 1,652 cross-view image sets from universities around the world. In recent months, several datasets [21, 22] have been proposed. However, all of them fail to address the issue of decentrality. To overcome this limitation, we propose CVSat, which provides a wider range of decentrality, facilitating a more in-depth investigation.
@@ 2.2 Cross-View Geo-Localization
To bridge the cross-view discrepancy, Shi et al. [13] first utilized polar transform on aerial-view images to achieve perspective transformation using geometry. This approach has since been widely adopted in subsequent research [23, 24, 25, 26]. Nevertheless, the initial application of the polar transform introduced significant distortions in the visual appearance. To counteract these distortions, Generative Adversarial Networks (GANs) have been employed, demonstrating efficacy in restoring the original appearance of transformed images [27, 28, 29, 30, 31]. With the introduction of the VIGOR [12], the polar transform-based methods are no longer applicable to non-center-aligned cross-view image pairs. Zhang et al. [15] proposed a feature recombination method that uses geometric spatial layout correspondence between cross-views to replace the polar transform. Though it has good performance on VIGOR, the design is not suitable for datasets with larger range of decentrality, like the proposed CVSat.
@@
Simultaneously, methods aimed at bridging the cross-view gap directly through the enhancement of feature representation [13, 32, 33, 34, 1, 35, 36, 37] are also in widely research. Several works use ViT as a backbone [24, 38, 22] for better feature extraction via encoding spatial information using self-attention mechanism. With the support of powerful backbone networks, recent feature-based methods [14] can achieve good performance, and some methods [39, 40] can even be trained in an unsupervised manner, albeit with limited performance.
3 CVSat dataset
@@ 3.1 Problem Statement
Given an area of interest (AOI), the goal of cross-view geo-localization is to determine the location of a street-view image arbitrarily within it by establishing correspondence with geo-tagged aerial-view images. Beyond previous datasets, the CVSat goes further in investigating the decentrality between cross-view images. As depicted in Figure 2(a), we define the area of the best matched street-view panorama of an aerial-view reference image as the hit area. Compared to VIGOR’s hit area (depicted as the yellow box), the hit area in CVSat (depicted as the red box) is notably larger, achieved through a lower overlap ratio of 12.5% (versus 50% in VIGOR). This disparity indicates a wider portion of decentrality among cross-view images in CVSat.
Figure 2: (a) Comparison of the hit area of VIGOR (yellow box) and CVSat (red box). Four subsets are divided within the hit area of CVSat with rising decentrality. (b) Comparison of the coverage scope between VIGOR and CVSat.
@@ 3.2 Decentrality Issue
In real-world applications of cross-view geo-localization, determining the overlap among reference images is crucial for constructing an effective reference image database, as it directly influences the database size, retrieval precision, and time consumption of the geo-localization process. However, this aspect has been understudied, and there is no unified industry standard. To address this gap, we conducted a comprehensive statistical analysis of various overlap levels among reference images, assessing their impact on the best-matched cross-view image pairs across all geo-locations within the CVSat dataset.
@@
Table 2: Statistical analysis of the reference dataset under varying overlap levels. S1-S4 represent the number of best-matched cross-view image pairs across four subsets, corresponding to different levels of decentrality. ”All” denotes the total number of images in the database, while ”Ratio” indicates the relative size compared to the baseline overlap of 12.5%.
We evaluated a range of overlap levels: 12.5%, 20%, 30%, 40%, and 50% among the reference aerial-view images. Maintaining the subset configurations from CVSat, as illustrated in Figure 2(a), we recorded the number of cross-view image pairs within subsets 1 to 4, each representing increasing levels of decentrality under different overlap levels. The results, detailed in Table 2, reveal that in 3 out of 5 cases of overlap settings, image pairs with large decentrality constitute a significant portion. For overlap levels of 12.5%, 20%, and 30%, image pairs in subset 4 (large decentrality) account for 13%, 31%, and 42%, respectively. This distribution pattern underscores that decentrality is a critical issue and needs to be fully investigated in CVGL.

### v2

@@ Abstract
Cross-View Geo-Localization tackles the challenge of image geo-localization in GNSS-denied environments, including disaster response scenarios, urban canyons, and dense forests, by matching street-view query images with geo-tagged aerial-view reference images. However, current research often relies on benchmarks and methods that assume center-aligned settings or account for only limited decentrality, which we define as the offset of the query image relative to the reference image center. Such assumptions fail to reflect real-world scenarios, where reference databases are typically pre-established without the possibility of ensuring perfect alignment for each query image. Moreover, decentrality is a critical factor warranting deeper investigation, as larger decentrality can substantially improve localization efficiency but comes at the cost of declines in localization accuracy. To address this limitation, we introduce DReSS (Decentrality Related Street-view and Satellite-view dataset), a novel dataset designed to evaluate cross-view geo-localization with a large geographic scope and diverse landscapes, emphasizing the decentrality issue. Meanwhile, we propose AuxGeo (Auxiliary Enhanced Geo-Localization) to further study the decentrality issue, which leverages a multi-metric optimization strategy with two novel modules: the Bird’s-eye view Intermediary Module (BIM) and the Position Constraint Module (PCM). BIM uses bird’s-eye view images derived from street-view panoramas as an intermediary, simplifying the “cross-view challenge with decentrality” to a cross-view problem and a decentrality problem. PCM leverages position priors between cross-view images to establish multi-grained alignment constraints. These modules improve the localization accuracy despite the decentrality problem. Extensive experiments demonstrate that AuxGeo outperforms previous methods on our proposed DReSS dataset, mitigating the issue of large decentrality, and also achieves state-of-the-art performance on existing public datasets such as CVUSA, CVACT, and VIGOR. The codes and dataset will be made available at https://github.com/SummerpanKing/DReSS.
@@ 1 Introduction
The cross-view coupling of street-view and satellite-view imagery in the spatial dimension is a prominent research focus in the field of remote sensing. This coupling not only facilitates better perspective observation of Earth’s surface (Li et al., 2023; Ye et al., 2024a) but also assists navigation in scenarios where GNSS signals are unavailable. Cross-View Geo-Localization (CVGL) addresses the challenge of image geo-localization by matching street-view query images with geo-tagged very high-resolution (VHR) satellite reference images from pre-established databases. It is particularly beneficial in complex environments where GNSS signals are unavailable, such as urban canyons with dense building clusters or heavily forested areas Ye et al. (2024b). This technology offers us an alternative way to localize ourselves in real scenarios, enabling emerging applications such as autonomous driving (Häne et al., 2017; Kim and Walter, 2017; Wan et al., 2016), robotic navigation (McManus et al., 2014), augmented reality (Chiu et al., 2018), geographic information aggregation (He et al., 2024), and disaster response (Li et al., 2024b).
@@
CVGL is challenging primarily due to the significant appearance changes between street-view and aerial-view images caused by viewpoint variations (Ling and Qin, 2022; Elhashash and Qin, 2022). Moreover, in real-world applications, it is not likely to have perfectly matched (center-aligned) reference images available for each query image, since the reference database is established with features generated by aerial images in advance. CVGL has to be robust and able to tolerate offsets between a street-view query and its corresponding aerial-view reference. Thus, we introduce the concept of decentrality, measuring the offset level between a query image position w.r.t. the center of its reference image. For the same search region, large decentrality requires less overlap between reference images, significantly reducing the number of reference images in the database and thereby improving the retrieval efficiency, which is essential for localization tasks. However, large decentrality presents substantial challenges to CVGL by diminishing content similarity between cross-view images and introducing disruptions caused by non-co-visible elements. Addressing these challenges requires robust partial matching mechanisms to effectively align cross-view images, as demonstrated in Figure 1. A detailed statistical analysis included in section 3.2 also highlights the importance of the decentrality issue in real-world applications.
@@
In existing datasets like CVUSA (Zhai et al., 2017) and CVACT (Liu and Li, 2019), the correspondences between cross-view images are assumed to be well center-aligned, which is over-simplified for real-world applications. The VIGOR dataset (Zhu et al., 2021) relaxes this restriction of center-aligned correspondence to some extent, while, a query image is still located within a quite small area around the center of its reference image, see Figure 2(a). While the VIGOR dataset defines some "semi-positive" street-view images with larger decentrality, there is always an alternative reference image available to form a "positive" pair for every street-view image. As a result, the VIGOR dataset is not well-suited for effectively investigating the decentrality issue, even if its original structure is reorganized. Therefore, we propose a novel dataset, DReSS, to evaluate CVGL in more realistic settings. DReSS distinguishes itself from previous datasets in two key ways. First, it allows for the evaluation of CVGL with larger decentrality (see Figure 2(a)) by sampling reference images seamlessly with a low overlap (12.5%). Second, DReSS comprises cross-view images collected from 8 cities worldwide, covering both urban and suburban areas with diverse landscapes, where the area of each reference image exceeds 400 $km^{2}$. This new large dataset enables a step forward in studying CVGL in a practical setting. A detailed comparison of decentrality among the above datasets is included in section 3.3.
@@
Large decentrality presents significant challenges for existing methods. Methods using polar transform, which convert aerial-view images into street-view-like images (Shi et al., 2019), are unsuitable without center-aligned settings. Other state-of-the-art methods (Deuser et al., 2023; Zhang and Zhu, 2024) do not take into consideration the impact of offsets between street-view and aerial-view images across a broader portion of decentrality, resulting in a significant decrease in retrieval precision. Therefore, we propose a novel method, AuxGeo (auxiliary enhanced geolocalization), to address the aforementioned problems. The AuxGeo employs multi-metric optimization with two innovative modules incorporated as auxiliary tasks to enhance the representation ability of the backbone. To mitigate viewpoint variation involving decentrality, we introduce the Bird’s-eye view Intermediary Module (BIM). The BIM uses BEV images derived from street-view panoramas as intermediaries, establishing additional connections between street-BEV and aerial-BEV images, simplifying the “cross-view challenge with decentrality” to a cross-view problem and a decentrality problem. To mitigate the problem of decentrality, we introduce the Position Constraint Module (PCM). The PCM leverages prior position correspondences between street-view and aerial-view images as supervision, establishing alignments in a multi-grained manner and enhancing the backbone’s ability to learn cross-view correspondences for feature representation even with significant decentrality.
@@
Extensive experiments demonstrate that AuxGeo improves retrieval precision, surpassing previous methods on our proposed DReSS dataset and existing public datasets. Our main contributions can be summarized as follows:
@@
  • We propose DReSS, the first public dataset designed to comprehensively evaluate CVGL methods in decentrality settings. It includes cross-view images from diverse geographic locations with extensive land coverage.
@@
  • Our proposed method AuxGeo outperforms previous approaches on the DReSS dataset, showing progressively greater improvements over the current methods as decentrality increases. It also achieves state-of-the-art localization accuracy on existing public datasets. Importantly, AuxGeo incurs no additional pre-processes or computational costs during inference.
The remainder of this paper is structured as follows: Section 2 provides a comprehensive review of related work, including existing datasets and geo-localization methods in the cross-view topic. Section 3 introduces the proposed DReSS dataset in detail and offers an in-depth analysis of the importance the decentrality issue. Section 4 outlines the proposed method, AuxGeo. Experimental results and discussions are presented in Sections 5 and 6, respectively. Finally, Section 7 concludes the paper with an analysis of the method’s advantages and limitations.
@@ 2.1 Cross-View Datasets
There are many datasets that have been introduced for cross-view geo-localization in the past decade (Lin et al., 2013; Vo and Hays, 2016; Tian et al., 2017). Among them, the most widely used datasets are CVUSA, CVACT, VIGOR and University-1652.
The original CVUSA is a huge dataset containing over 1 million ground and aerial images from multiple cities in the United States, first proposed by Workman et al. (Workman et al., 2015). The current used version of CVUSA is a subset made by Zhai et al. (Zhai et al., 2017), including 35,532 street-aerial image pairs for training and 8,884 pairs for testing. Similar to CVUSA, Liu et al. (Liu and Li, 2019) proposed CVACT which has 35,532 image pairs for training, 8,884 pairs for validation and 92,802 pairs for testing. Extending beyond conventional one-to-one retrieval, Zhu et al. (Zhu et al., 2021) introduced VIGOR. This dataset comprises 105,214 street-view images and 90,618 aerial-view images, uniquely characterized by its assumption of random placements within the target area without center-aligned settings. VIGOR is regarded as the most realistic dataset for cross-view geo-localization, as it extends the relationships between cross-view images beyond merely center-aligned correspondence. However, even with this advancement, the decentrality of positive samples in VIGOR is still confined to a relatively centered area, which restricts its potential for a more thorough exploration of the issue. University-1652 (Zheng et al., 2020) is the first dataset to simultaneously include images from satellite, synthetic drone, and ground levels, providing 1,652 cross-view image sets from universities around the world. In recent months, several datasets (Ye et al., 2025; Huang et al., 2024) have been proposed. However, all of them fail to address the issue of decentrality. To overcome this limitation, we propose DReSS, which provides a wider range of decentrality, facilitating a more in-depth investigation.
@@ 2.2 Cross-View Geo-Localization
To bridge the cross-view discrepancy, Shi et al. (Shi et al., 2019) first utilized polar transform on aerial-view images to achieve perspective transformation using geometry. This approach has since been widely adopted in subsequent research (Shi et al., 2020a; Yang et al., 2021; Wang et al., 2023; Zhang et al., 2023). Nevertheless, the initial application of the polar transform introduced significant distortions in the visual appearance. To counteract these distortions, Generative Adversarial Networks (GANs) have been employed, demonstrating efficacy in restoring the original appearance of transformed images (Toker et al., 2021; Regmi and Borji, 2018; Regmi and Shah, 2019; Lu et al., 2020; Shi et al., 2022). With the introduction of the VIGOR (Zhu et al., 2021), the polar transform-based methods are no longer applicable to non-center-aligned cross-view image pairs. Zhang et al. (Zhang and Zhu, 2024) proposed a feature recombination method that uses geometric spatial layout correspondence between cross-views to replace the polar transform. Though it has good performance on VIGOR, the design is not suitable for datasets with larger range of decentrality, like the proposed DReSS.
@@
Simultaneously, methods aimed at bridging the cross-view gap directly through the enhancement of feature representation (Shi et al., 2019, 2020b; Hu et al., 2018; Sun et al., 2019; Ye et al., 2024b; Cheng et al., 2018; Wu et al., 2024; Xia et al., 2024) are also in widely research. Several works use ViT as a backbone (Yang et al., 2021; Zhu et al., 2022; Huang et al., 2024) for better feature extraction via encoding spatial information using self-attention mechanism. With the support of powerful backbone networks, recent feature-based methods (Deuser et al., 2023) can achieve good performance, and some methods (Li et al., 2024a, c) can even be trained in an unsupervised manner, albeit with limited performance.
3 DReSS dataset
@@ 3.1 Problem Statement
Given an area of interest (AOI), the goal of cross-view geo-localization is to determine the location of a street-view image arbitrarily within it by establishing correspondence with geo-tagged aerial-view images. Beyond previous datasets, the DReSS goes further in investigating the decentrality between cross-view images. As depicted in Figure 2(a), we define the area of the best matched street-view panorama of an aerial-view reference image as the hit area. Compared to VIGOR’s hit area (depicted as the yellow box), the hit area in DReSS (depicted as the red box) is notably larger, achieved through a lower overlap ratio of 12.5% (versus 50% in VIGOR). This disparity indicates a wider portion of decentrality among cross-view images in DReSS.
Figure 2: (a) Comparison of the hit area of VIGOR (yellow box) and DReSS (red box). Four subsets are divided within the hit area of DReSS with rising decentrality. (b) Comparison of the coverage scope between VIGOR and DReSS.
@@ 3.2 Decentrality Issue
In real-world applications of cross-view geo-localization, determining the overlap among reference images is crucial for constructing an effective reference image database, as it directly influences the database size, retrieval precision, and time consumption of the geo-localization process. However, this aspect has been understudied, and there is no unified industry standard. To address this gap, we conducted a comprehensive statistical analysis of various overlap levels among reference databases, assessing their impact on the best-matched cross-view image pairs across all geo-locations within the DReSS dataset.
@@
Table 2: Statistical analysis of the reference dataset under varying overlap levels. S1-S4 represent the number of best-matched cross-view image pairs across four subsets, corresponding to different levels of decentrality. "All" denotes the total number of images in the database, while "Ratio" indicates the relative size compared to the baseline overlap of 12.5%.
We evaluated a range of overlap levels: 12.5%, 20%, 30%, 40%, and 50% among the reference aerial-view images. Maintaining the subset configurations from DReSS, as illustrated in Figure 2(a), we recorded the number of cross-view image pairs within subsets 1 to 4, each representing increasing levels of decentrality under different overlap levels. The results, detailed in Table 2, reveal that in 3 out of 5 cases of overlap settings, image pairs with large decentrality constitute a significant portion. For overlap levels of 12.5%, 20%, and 30%, image pairs in subset 4 (large decentrality) account for 13%, 31%, and 42%, respectively. This distribution pattern underscores that decentrality is a critical issue and needs to be fully investigated in CVGL.